# Mechanistic Interpretability of Large-Scale Counting in LLMs through a System-2 Strategy

### v1

@@ Abstract
Large language models (LLMs), despite strong performance on complex mathematical problems, exhibit systematic limitations in counting tasks. This issue arises from architectural limits of transformers, where counting is performed across layers, leading to degraded precision for larger counting problems due to depth constraints. To address this limitation, we propose a simple test-time strategy inspired by System-2 cognitive processes that decomposes large counting tasks into smaller, independent sub-problems that the model can reliably solve. We evaluate this approach using observational and causal mediation analyses to understand the underlying mechanism of this System-2-like strategy. Our mechanistic analysis identifies key components: latent counts are computed and stored in the final item representations of each part, transferred to intermediate steps via dedicated attention heads, and aggregated in the final stage to produce the total count. Experimental results demonstrate that this strategy enables LLMs to surpass architectural limitations and achieve high accuracy on large-scale counting tasks. This work provides mechanistic insight into System-2 counting in LLMs and presents a generalizable approach for improving and understanding their reasoning behavior.
@@ 1 Introduction
Counting is a fundamental cognitive operation that underpins a wide range of reasoning tasks, from basic arithmetic to more complex forms of quantitative analysis (feigenson2004core; dehaene2011numbersense). In Large Language Models (LLMs), the ability to count is important for controlled generation such as length-constrained summarization, sequential enumeration, and broader numerical/arithmetic reasoning (retkowski-waibel-2025-zero; hou2025sequentialenumeration; yang2024numbercookbook; gambardella-etal-2024-language).
Recent research has provided insights into the counting mechanism of LLMs, demonstrating that these models use a layerwise internal counting process where numerical information is progressively accumulated across transformer layers (hasani2025countscope). However, due to the depth limits, this progressive counting mechanism becomes saturated as the number of items increases. Complementary work on numerical representations suggests that LLMs encode numbers in a compressed, sublinear manner, similar to the human logarithmic mental number line (alquboj2025numberrepresentations; dehaene2011numbersense). While numerical order is preserved, representational resolution decreases with magnitude, explaining the reduced precision for large values.
@@
These findings suggest that LLMs struggle to accurately count large numbers of items, with performance typically degrading for two- and three-digit counts (zhang2024countingabilitytokenization; yehudai2024whencantransformerscount; fu2024lettercounting). This limitation reflects a fundamental architectural constraint rather than a lack of training data or supervision, as more layers are required to count a larger number of items (vaswani2017attention; yehudai2024whencantransformerscount; golkar2024contextualcounting). We argue that this failure stems from the model’s reliance on a System-1–like processing approach, which is fast, automatic, and capacity-limited (kahneman2011thinking; dehaene2011numbersense).
To address this issue, we propose a simple test-time strategy that adopts a System-2–like approach to counting (Figure 1(kahneman2011thinking). Instead of relying on the model’s internal counting mechanisms, which are constrained by architectural limits, our approach decomposes large counting tasks into smaller, independent sub-tasks (radhakrishnan2023questiondecomposition; qharabagh2024lvlmcount). Each sub-task, containing a manageable number of items, can be reliably counted by the model. The results from each sub-task are then aggregated to produce the final count. This approach mirrors human cognitive strategies, where complex problems are broken down into simpler, easier-to-solve sub-problems (kahneman2011thinking; dehaene2011numbersense).
Our behavioral experiments on various LLMs demonstrate the effectiveness of this strategy in overcoming architectural limitations without requiring model modifications or fine-tuning. Additionally, we provide a detailed mechanistic interpretation of how it functions within LLMs. Using attention analysis and causal mediation techniques (heimersheim2024activationpatching; geiger2021causalabstractions; ghandeharioun2024patchscopes; zhang2024activationpatching_bestpractices), we trace the flow of numerical information across the model and identify the mechanisms mediating the System-2 counting process. Figure 2 provides an overview of the main components of the System-2 counting mechanism and its internal information flow. This work offers a new perspective on both improving and understanding LLMs’ reasoning capabilities, with a focus on a fundamental cognitive task.
@@ 2 Problem Setup and Methodology
Qwen2.5 7B Unstructured w/o steps 0.38 0.13 0.06 0.00 0.88 2.19 5.29 10.50
@@
Llama 3 8B Unstructured w/o steps 0.80 0.49 0.02 0.00 0.20 0.65 4.93 11.92
@@
Gemma 3 27B Unstructured w/o steps 1.00 0.70 0.40 0.00 0.00 0.30 1.00 4.90
@@
Table 1: Average accuracy and mean absolute error (MAE) across open-source models for context sizes from 11 to 50. Each model is evaluated on structured and unstructured inputs, with and without intermediate reasoning steps.
@@
We follow the standard counting framework used in prior research (hasani2025countscope). In this setup, a list of repeated items, such as fruits or animals, is presented to the model, and the task is to report the total number of items. For example, given a context like “apple, apple, apple, …”, the model must output the total number of items in the list. Previous work has shown that LLMs exhibit high accuracy for small counts (fewer than 10 items), but that performance deteriorates as the number of items increases beyond this range (zhang2024countingabilitytokenization; fu2024lettercounting). This suggests that the model’s counting ability is limited by the depth of the transformer architecture and its internal counting mechanism (yehudai2024whencantransformerscount).
To overcome this limitation, motivated by prior work on partitioning images in visual reasoning tasks (izadi2025viser), we introduce a simple strategy that explicitly partitions the input list into smaller sub-problems. We use an external separator (|) to divide the list into smaller partitions. For instance, a structured context with three partitions is: “apple, apple, apple, apple, apple | apple, apple, apple | apple, apple, apple, apple, apple, apple”.
The model is then instructed to first count the items in each partition and to aggregate the partial counts to produce the final result. This ensures that each sub-problem remains within the model’s reliable counting regime. The number of items in each partition is chosen randomly from a range that the model can handle accurately. Models with deeper architectures can reliably process larger partitions.
Based on the input structure and output format, we consider four baselines in our study. Two input formats are used: an unstructured context with comma-separated items and a structured context with partitions separated by vertical bar symbol (|). For generation, we evaluate two output variants: a short-answer format, where only the final count is produced (corresponding to an immediate System-1-like process), and a Chain-of-Thought (CoT) format, where intermediate reasoning steps are included before the final answer (corresponding to a System-2-like process).
3 Behavioral Results
Figure 3: Decoded output probabilities for the unstructured baseline method on Qwen2.5 7B. The heatmap shows the decoded probabilities of model outputs, averaged over different item types, for target counts ranging from 1 to 25. As the count increases beyond 10, the diagonal entries gradually fade, indicating reduced model confidence.
To illustrate how LLMs lose counting precision in long contexts, we first measure the average prediction probability of Qwen2.5 7B (qwen25) across different context sizes. As shown in Figure 3, model confidence decreases as the number of items exceeds 10. We also observe systematic biases toward more frequent numbers. For example, target counts of 16 and 21 are often predicted as 15 and 20.
@@ 3 Behavioral Results
We then conduct more systematic experiments on both open-source and closed-source models. For open-source models, we evaluate Qwen2.5 7B, Llama 3 8B (llama3), and Gemma 3 27B (gemma), which have 28, 32, and 62 layers, respectively. In addition, we evaluate GPT-4o (Achiam2023GPT4) and Gemini-2.5-Pro (gemini2025pushing) as stronger proprietary models on longer contexts. The corresponding results are reported in Tables 1 and 2. Across all models, we find that only one configuration consistently succeeds on long contexts: structured inputs combined with intermediate reasoning steps.
@@
Surprisingly, encouraging CoT reasoning alone, without structured input, does not yield a notable improvement. Furthermore, structured input without intermediate steps is also ineffective and can be harmful in some cases. A likely explanation is that models must first consolidate partial results in intermediate steps before aggregating them into the final answer through a two-stage process. Our observational and causal analyses provide further insight into why models cannot simultaneously gather partial counts from the context and add them up without loss. Overall, these results show that neither external structure nor reasoning alone is sufficient. Their combination is necessary to overcome large-scale counting failures.
@@ 4 Attention Analysis
We analyze attention patterns of Qwen2.5 7B (qwen25) to understand how the proposed System-2 strategy enables large-scale counting. We first identify which tokens are attended to when the model generates (i) intermediate partition-level counts and (ii) the final aggregated answer. We then localize the layers and heads that contribute most to these processes.
Figure 4(a) shows attention patterns to the input items when the model generates intermediate reasoning steps. We consider a structured context with three partitions containing 3, 4, and 5 items. During the generation of a reasoning token such as “part 2: 4”, the attention weight of the generated number peaks sharply on the final item and the final comma separator of the corresponding partition. This pattern is consistent across prompts and is most pronounced in middle to late layers, typically around layers 19 to 23. Earlier items within the same partition receive substantially lower attention.
@@
To further localize the attention pathways involved in intermediate reasoning and final aggregation, we focus on the most relevant tokens identified in Figure 4. We average attention values across different configurations, including item types and partition sizes. Figure 5 shows that attention peaks in layers 19 to 23, highlighting the central role of these layers in information transfer and aggregation. Figure 6 further shows that the most influential heads are concentrated in layers 21, 22, and 23. For example, head 13 in layer 22 consistently exhibits high attention to the selected tokens.
@@ 5 Causal Mediation Analysis
To assess the hypothesis of a multi-stage System-2 counting mechanism suggested by the attention analysis, we perform a set of causal mediation experiments based on activation patching (heimersheim2024activationpatching; zhang2024activationpatching_bestpractices), masking ablation, and attention knockout (attention_knockout). The goal is to identify where partition-level count information is stored, how it is transferred and consolidated, and which components are causally required for correct aggregation.
@@ 5.1 Token-Level Information Probing
We first examine where partition-level counts are encoded in the input context. This analysis requires vocabulary projection and probing tools. We observe that existing tools, such as logit-lens (nostalgebraist2020logitlens) and tuned-lens (belrose2023tunedlens), are not reliable for decoding numerical information. We therefore use the CountScope method (hasani2025countscope) to decode the implicit count associated with tokens. CountScope is a causal probing method that patches the activation of a target token into a blank counting context and decodes the model’s output as the implied count.
@@ 5.3 Information Pathway Localization
To perform fine-grained circuit localization, we analyze the pathways responsible for System-2 counting using attention knockout (attention_knockout). We selectively block individual attention heads and layers and measure the resulting drop in counting accuracy for intermediate and final steps. Figure 9 shows that for Qwen2.5 7B, the most influential components are concentrated in layer 22. In particular, attention head 13 is most important for intermediate steps, while head 1 is most important for the final aggregation step.
@@ 5.4 Cross-Context Activation Patching
Finally, we perform cross-context activation patching to test how partition-level information is combined to form the final answer. We focus on intermediate responses and final aggregation steps. Our attention and masking analyses indicate that transferred numerical information from the context to intermediate steps is consolidated in the specific tokens “:”, whitespace, and the partition number. Here, we investigate this behavior using a causal intervention setup (geiger2021causalabstractions).
@@ 6 Related Work
A growing body of work shows that counting in LLMs is brittle: performance can hinge on surface form and segmentation rather than a stable procedure (fu2024lettercounting). Subword tokenization can blur item boundaries and measurably affect counting accuracy (zhang2024countingabilitytokenization). More broadly, deterministic-task evaluations (including counting) can flip under minor prompt/content changes, complicating extrapolation from narrow setups (ball2024fixed_effect_fallacy). On the theory side, transformers can exactly count token frequencies only in specific regimes, with feasibility tied to representation capacity and context-length scaling (yehudai2024whencantransformerscount).
@@
Cognitive work argues for specialized number systems and structured magnitude representations (feigenson2004core; dehaene2011numbersense), motivating analogous questions in neural models. Representation analyses study how numerical magnitude is organized in LLM hidden spaces and relate these structures to human-like effects (alquboj2025numberrepresentations). Mechanistic studies localize how counting signals emerge and update across items in controlled settings (golkar2024contextualcounting). Closest to our setting, hasani2025countscope combine a target-context probe with layerwise/token-level analyses and causal interventions to identify latent counter-like signals across both LLMs and LVLMs.
@@
Chain-of-Thought can improve multi-step reasoning by eliciting intermediate computations (wei2022cot), but explicit traces are not necessarily faithful. Question decomposition work studies when decomposition increases faithfulness versus mainly serving as scaffolding (radhakrishnan2023questiondecomposition).
Counting in LVLMs.
Counting remains challenging for LVLMs, particularly under clutter, layout variation, and compositional settings that stress binding between category and quantity (guo2025vlmcountbench; campbell2024binding). Diagnostics and targeted modifications analyze where LVLM counting breaks and how performance can be recovered (alghisi2025dereconstructing), while structured visual inputs or supervision can improve counting (izadi2025viser; qharabagh2024lvlmcount; grounding_id). Mechanistic evidence further suggests layerwise latent counter signals distributed across visual tokens (hasani2025countscope).
@@
Causal interventions such as activation patching help identify computation-critical states (heimersheim2024activationpatching; geiger2021causalabstractions), with best-practice guidance for reliable metrics and methods (zhang2024activationpatching_bestpractices). Patch-based inspection frameworks further systematize intervention/inspection configurations (ghandeharioun2024patchscopes). Readout-style probes provide a complementary perspective by tracking when information becomes linearly decodable across layers, including the Logit Lens (nostalgebraist2020logitlens) and learned variants such as the Tuned Lens (belrose2023tunedlens).
Our work.
Our setup is closest to System-2-style prompting via intermediate steps: like CoT, we elicit explicit intermediate computations (wei2022cot), and like task decomposition, we solve the instance by breaking it into sequential subproblems (radhakrishnan2023questiondecomposition). The key difference is that the decomposition is tied to a structured input format: we first partition the input into marked segments and then use a CoT protocol that is aligned with this structure (produce segment-level subcounts, then aggregate). This creates explicit stage boundaries that are absent in monolithic counting prompts and lets us study how counting representations evolve across stages. Mechanistically, our analysis aligns with prior work that localizes latent counter-like signals and tests them with interventions (hasani2025countscope), while shifting the question to how the structured CoT decomposition reshapes internal computation on long contexts. In this way, we connect behavioral brittleness observations (ball2024fixed_effect_fallacy; zhang2024countingabilitytokenization) to a stage-wise mechanistic account of how count information is formed, routed, and combined under structured prompting (heimersheim2024activationpatching; zhang2024activationpatching_bestpractices).

### v2

@@ Abstract
Large language models (LLMs), despite strong performance on complex mathematical problems, exhibit systematic limitations in counting tasks. This issue arises from the architectural limits of transformers, where counting is performed across layers, leading to degraded precision for larger counting problems due to depth constraints. To address this limitation, we propose a simple test-time strategy inspired by System-2 cognitive processes that decomposes large counting tasks into smaller, independent sub-problems that the model can reliably solve. We evaluate this approach using observational and causal mediation analyses to understand the underlying mechanism of this System-2-like strategy. Our mechanistic analysis identifies key components: latent counts are computed and stored in the final item representations of each part, transferred to intermediate steps via dedicated attention heads, and aggregated in the final stage to produce the total count. Experimental results demonstrate that this strategy enables LLMs to surpass architectural limitations and achieve higher accuracy on large-scale counting tasks. This work provides mechanistic insight into System-2 counting in LLMs and presents a generalizable approach for improving and understanding their reasoning behavior.
@@ 1 Introduction
Counting is a fundamental cognitive operation that underpins a wide range of reasoning tasks, from basic arithmetic to more complex forms of quantitative analysis (Feigenson et al., 2004; Dehaene, 2011). In large language models (LLMs), the ability to count is important for controlled generation such as length-constrained summarization, sequential enumeration, and broader numerical/arithmetic reasoning (Retkowski and Waibel, 2025; Hou et al., 2025; Yang et al., 2024c; Gambardella et al., 2024).
Recent research has provided insights into the counting mechanism of LLMs, demonstrating that these models use a layerwise internal counting process where numerical information is progressively accumulated across transformer layers (Hasani et al., 2025b). However, due to the depth limits, this progressive counting mechanism becomes saturated as the number of items increases. Complementary work on numerical representations suggests that LLMs encode numbers in a compressed, sublinear manner, similar to the human logarithmic mental number line (AlquBoj et al., 2025; Dehaene, 2011). While numerical order is preserved, representational resolution decreases with magnitude, explaining the reduced precision for large values.
@@
These findings suggest that LLMs struggle to accurately count large numbers of items, with performance typically degrading for two- and three-digit counts (Zhang et al., 2024; Yehudai et al., 2024; Fu et al., 2024). This limitation reflects a fundamental architectural constraint rather than a lack of training data or supervision, as more layers are required to count a larger number of items (Vaswani et al., 2017; Yehudai et al., 2024; Golkar et al., 2024). We argue that this failure stems from the model’s reliance on a System-1–like processing approach, which is fast, automatic, and capacity-limited (Kahneman, 2011; Dehaene, 2011).
To address this issue, we propose a simple test-time strategy that adopts a System-2–like approach to counting (Figure 1(Kahneman, 2011). Instead of relying on the model’s internal counting mechanisms, which are constrained by architectural limits, our approach decomposes large counting tasks into smaller, independent sub-tasks (Radhakrishnan et al., 2023; Qharabagh et al., 2024). Each sub-task, containing a manageable number of items, can be reliably counted by the model. The results from each sub-task are then aggregated to produce the final count. This approach mirrors human cognitive strategies, where complex problems are broken down into simpler, easier-to-solve sub-problems (Dehaene, 2011). In this study, we use System-1 and System-2 as high-level operational abstractions inspired by human cognition (Kahneman, 2011). Specifically, we refer to System-1 as the model’s single-pass, layer-wise processing, and to System-2 as a more deliberate procedure that extends computation across token generation beyond a fixed layer-wise computation.
Our behavioral experiments on various LLMs demonstrate the effectiveness of this strategy in overcoming architectural limitations without requiring model modifications or fine-tuning. Additionally, we provide a detailed mechanistic interpretation of how it functions within LLMs. Using attention analysis and causal mediation techniques (Heimersheim and Nanda, 2024; Geiger et al., 2021; Ghandeharioun et al., 2024; Zhang and Nanda, 2024), we trace the flow of numerical information across the model and identify the mechanisms mediating the System-2 counting process. Figure 2 provides an overview of the main components of the System-2 counting mechanism and its internal information flow. This work offers a new perspective on both improving and understanding LLMs’ reasoning capabilities, with a focus on a fundamental cognitive task.
@@ 2 Problem Setup and Methodology
We follow the standard counting framework used in prior research (Hasani et al., 2025b). In this setup, a list of repeated items, such as fruits or animals, is presented to the model, and the task is to report the total number of items. For example, given a context like “apple, apple, apple, …”, the model must output the total number of items in the list. Previous work has shown that LLMs exhibit high accuracy for small counts (fewer than 10 items), but that performance deteriorates as the number of items increases beyond this range (Zhang et al., 2024; Fu et al., 2024). This suggests that the model’s counting ability is limited by the depth of the transformer architecture and its internal counting mechanism (Yehudai et al., 2024).
To overcome this limitation, motivated by prior work on partitioning images in visual reasoning tasks (Izadi et al., 2025), we introduce a simple strategy that explicitly partitions the input list into smaller sub-problems. We use an external separator (|) to divide the list into smaller partitions. For instance, a structured context with three partitions is: “apple, apple, apple, apple, apple | apple, apple, apple | apple, apple, apple, apple, apple, apple”.
The model is then instructed to first count the items in each partition and to aggregate the partial counts to produce the final result. This ensures that each sub-problem remains within the model’s reliable counting regime. The number of items in each partition is chosen randomly from a range that the model can handle accurately. Models with deeper architectures can reliably process larger partitions. This strategy is based on test-time scaling by leveraging the LLM’s inherent capabilities and does not require fine-tuning or any external tools.
Based on the input structure and output format, we consider four baselines in our study. Two input formats are used: an unstructured context with comma-separated items and a structured context with partitions separated by vertical bar symbol (|). For generation, we evaluate two output variants: a short-answer format, where only the final count is produced (corresponding to an immediate System-1-like process), and a Chain-of-Thought (CoT) (Wei et al., 2022) format, where intermediate reasoning steps are included before the final answer (corresponding to a System-2-like process). Full details of the input formats, partitioning, and prompting strategies are provided in Appendix A.1.
3 Behavioral Results
Figure 3: Decoded output probabilities for the unstructured baseline method on Qwen2.5-7B. The heatmap shows the decoded probabilities of model outputs, averaged over different item types, for target counts ranging from 1 to 25. As the count increases beyond 10, the diagonal entries gradually fade, indicating reduced model confidence.
To illustrate how LLMs lose counting precision in long contexts, we first measure the average prediction probability of Qwen2.5-7B (Yang et al., 2024a) across different context sizes. As shown in Figure 3, model confidence decreases as the number of items exceeds 10. We also observe systematic biases toward more frequent numbers. For example, target counts of 16 and 24 are often predicted as 15 or 25, with 21 and 22 typically predicted as 20.
@@ 3 Behavioral Results
Qwen2.5-7B Unstructured w/o steps 0.38 0.13 0.06 0.00 0.88 2.19 5.29 10.50
@@
Llama3-8B Unstructured w/o steps 0.80 0.49 0.02 0.00 0.20 0.65 4.93 11.92
@@
Gemma3-27B Unstructured w/o steps 1.00 0.70 0.40 0.00 0.00 0.30 1.00 4.90
@@
Table 1: Average accuracy and mean absolute error (MAE) across open-source models for context sizes from 11 to 50. Each model is evaluated on structured and unstructured inputs, with and without intermediate reasoning steps. MAE complements exact-match accuracy by measuring how far incorrect predictions are from the true count.
@@
We then conduct more systematic experiments on both open-source and closed-source models, covering different model sizes, depths, training strategies, and tokenization schemes. For open-source models, we evaluate Qwen2.5-7B, Llama3-8B (Grattafiori et al., 2024), and Gemma3-27B (Kamath et al., 2025), which have 28, 32, and 62 layers, respectively. In addition, we evaluate GPT-4o (Achiam et al., 2023) and Gemini-2.5-Pro (Team, 2025) as stronger proprietary models on longer contexts  [^1] For these stronger closed-source models, performance is near-perfect for shorter contexts (below 50 items). We therefore use longer contexts, where the task is no longer saturated and differences between strategies become visible.. The corresponding results are reported in Tables 1 and 2. Overall, performance decreases as context size increases and is associated with model scale, especially depth, with larger and deeper models showing higher accuracy. For each model, we find that only one configuration consistently succeeds on long contexts: structured inputs combined with intermediate reasoning steps.
@@
Surprisingly, encouraging CoT reasoning alone, without structured input, does not yield a notable improvement. Furthermore, structured input without intermediate steps is also ineffective and can be harmful in some cases. These results show that neither external structure nor reasoning alone is sufficient. Their combination is necessary to overcome large-scale counting failures. In many models, the unstructured short-answer baseline performs better than the structured short-answer baseline. This suggests that the model must first consolidate partial results in intermediate steps before aggregating them into the final answer through a two-stage process.
A likely explanation for the failure of the structured short-answer setting is related to the latent counting mechanism. Prior work showed that some models tend to output the maximum latent count observed in the context rather than the true total count (Hasani et al., 2025b). In our structured format, the counter resets after each partition, so the maximum latent count often matches the largest partition size instead of the total sum. Consistent with this view, for context sizes 11 to 30, 13% of Qwen2.5-7B errors and 43.6% of Llama3-8B errors exactly match the largest partition size. In addition, the high MAE of this setting in Table 1 suggests systematic bias in erroneous predictions. Our observational and causal analyses provide further insight into why models cannot simultaneously gather partial counts from the context and add them up without loss.
Finally, we analyze failure cases of the structured CoT setting by separating errors into intermediate partition counts and final aggregation. Table 3 shows that the main source of failure is the intermediate steps. This indicates that once correct partition counts are produced, most models can reliably add them.
Model Total Acc Final Step Acc Interm. Acc
Qwen2.5 7B 0.51 0.86 0.53
Llama 3 8B 0.49 0.96 0.48
Gemma 3 27B 0.71 0.93 0.76
GPT-4o 0.89 1.00 0.89
Gemini-2.5-Pro 0.94 0.97 0.94
Table 3: Failure analysis of the structured CoT setting. Total accuracy measures exact final answers. Final-step accuracy measures correctness of the aggregation step conditioned on generated intermediate counts. Intermediate-step accuracy measures correctness of partition-level counts.
@@
Additional experiments on smaller and math-specialized models are presented in Appendix A.2, showing that the observed trends generalize across model scale and domain specialization. Robustness checks across tokenizers and alternative input structures are provided in Appendix A.3, and the overall trends hold across tokenizers and input formats, demonstrating that the System-2 mechanism is robust and generalizable.
@@ 4 Attention Analysis
We analyze attention patterns of Qwen2.5-7B (Yang et al., 2024a) to understand how the proposed System-2 strategy enables large-scale counting. We first identify which tokens are attended to when the model generates (i) intermediate partition-level counts and (ii) the final aggregated answer. We then localize the layers and heads that contribute most to these processes.
Figure 4(a) shows attention patterns to the input items when the model generates intermediate reasoning steps. We consider a structured context with three partitions containing 3, 4, and 6 items. During the generation of a reasoning token such as “part 2: 4”, the attention weight of the generated number peaks sharply on the final item and the final comma separator of the corresponding partition. This pattern is consistent across prompts and is most pronounced in middle to late layers, typically around layers 19 to 23. Earlier items within the same partition receive substantially lower attention.
@@
To further localize the attention pathways involved in intermediate reasoning and final aggregation, we focus on the most relevant tokens identified in Figure 4. We average attention values across different configurations, including item types and partition sizes. Figure 5 shows that attention peaks in layers 19 to 23, highlighting the central role of these layers in information transfer and aggregation. In addition, Figure 6 further shows that the most influential heads are concentrated in layers 21, 22, and 23. For example, head 13 in layer 22 consistently exhibits high attention to the selected tokens. Additional methodological details and attention analyses across multiple model architectures are provided in Appendix A.4.
@@ 5 Causal Mediation Analysis
To assess the hypothesis of a multi-stage System-2 counting mechanism suggested by the attention analysis, we perform a set of causal mediation experiments based on activation patching (Heimersheim and Nanda, 2024; Zhang and Nanda, 2024), masking ablation, and attention knockout (Geva et al., 2023). The goal is to identify where partition-level count information is stored, how it is transferred and consolidated, and which components are causally required for correct aggregation.
@@ 5.1 Token-Level Information Probing
We first examine where partition-level counts are encoded in the input context. This analysis requires vocabulary projection and probing tools. We observe that existing tools, such as logit-lens (NostAlgebraist, 2020) and tuned-lens (Belrose et al., 2023), are not reliable for decoding numerical information. We therefore use the CountScope method (Hasani et al., 2025b) to decode the implicit count associated with tokens. CountScope is a causal probing method that patches the activation of a target token into a blank counting context and decodes the model’s output as the implied count.
@@ 5.2 Token-Level Causal Interventions
(a) Layer masking (deletion)
(b) Layer unmasking (insertion)
Figure 9: Average change in log probability of intermediate counts after layer masking (deletion) and unmasking (insertion) of the final item and separator of each partition. In the masking experiment, embeddings from the target layer are zero-ablated. In the unmasking experiment, embeddings of the selected tokens are zero-ablated across all layers, and only the target-layer embeddings are restored from the clean run.
@@ 5.3 Information Pathway Localization
To perform fine-grained circuit localization, we analyze the pathways responsible for System-2 counting using attention knockout (Geva et al., 2023). We selectively block individual attention heads and layers and measure the resulting drop in counting accuracy for intermediate and final steps. Figure 9 shows that for Qwen2.5-7B, the most influential components are concentrated in layer 22. In particular, attention head 13 is most important for intermediate steps, while head 1 is most important for the final aggregation step (Figure 10).
@@ 5.4 Cross-Context Activation Patching
Finally, we perform cross-context activation patching to test how partition-level information is combined to form the final answer. We focus on intermediate responses and final aggregation steps. Our attention and masking analyses indicate that transferred numerical information from the context to intermediate steps is consolidated in the specific tokens “:”, the whitespace token, and the partition number. Here, we investigate this behavior using a causal intervention setup (Geiger et al., 2021).
@@ 6 Related Work
A growing body of work shows that counting in LLMs is brittle: performance can hinge on surface form and segmentation rather than a stable procedure (Fu et al., 2024). Subword tokenization can blur item boundaries and measurably affect counting accuracy (Zhang et al., 2024). More broadly, deterministic-task evaluations (including counting) can flip under minor prompt/content changes, complicating extrapolation from narrow setups (Ball et al., 2024). On the theory side, transformers can exactly count token frequencies only in specific regimes, with feasibility tied to representation capacity and context-length scaling (Yehudai et al., 2024).
@@
Cognitive work argues for specialized number systems and structured magnitude representations (Feigenson et al., 2004; Dehaene, 2011), motivating analogous questions in neural models. Representation analyses study how numerical magnitude is organized in LLM hidden spaces and relate these structures to human-like effects (AlquBoj et al., 2025). Mechanistic studies localize how counting signals emerge and update across items in controlled settings (Golkar et al., 2024). Closest to our setting, Hasani et al. (2025b) combine a target-context probe with layerwise/token-level analyses and causal interventions to identify latent counter-like signals across both LLMs and LVLMs.
@@
Chain-of-Thought can improve multi-step reasoning by eliciting intermediate computations (Wei et al., 2022), but explicit traces are not necessarily faithful. Question decomposition work studies when decomposition improves faithfulness rather than mainly serving as scaffolding (Radhakrishnan et al., 2023). Decomposing images into smaller regions has also been shown to be effective for visual reasoning (Izadi et al., 2025; Hasani et al., 2025a). This strategy improves visual counting performance (Izadi et al., 2025; Qharabagh et al., 2024; Hasani et al., 2025a), but its underlying mechanism remains underexplored.
@@
Causal interventions such as activation patching help identify computation-critical states (Heimersheim and Nanda, 2024; Geiger et al., 2021), with best-practice guidance for reliable metrics and methods (Zhang and Nanda, 2024). Patch-based inspection frameworks further systematize intervention/inspection configurations (Ghandeharioun et al., 2024). Readout-style probes provide a complementary perspective by tracking when information becomes linearly decodable across layers, including the Logit Lens (NostAlgebraist, 2020) and learned variants such as the Tuned Lens (Belrose et al., 2023).
Positioning of This Work.
Our setup is closest to System-2-style prompting via intermediate steps: like CoT, we elicit explicit intermediate computations (Wei et al., 2022), and like task decomposition, we solve the instance by breaking it into sequential subproblems (Radhakrishnan et al., 2023). The key difference is that the decomposition is tied to a structured input format: following prior vision-based studies (Izadi et al., 2025; Hasani et al., 2025a), we first partition the input text into marked segments and then use a CoT protocol aligned with this structure by producing segment-level subcounts and then aggregating them. This creates explicit stage boundaries that are absent in monolithic counting prompts and lets us study how counting representations evolve across stages. Mechanistically, our analysis aligns with prior work that localizes latent counter-like signals and tests them with interventions (Hasani et al., 2025b), while shifting the question to how the structured CoT decomposition reshapes internal computation on long contexts. In this way, we connect behavioral brittleness observations (Ball et al., 2024; Zhang et al., 2024) to a stage-wise mechanistic account of how count information is formed, routed, and combined under structured prompting (Heimersheim and Nanda, 2024; Zhang and Nanda, 2024).